# Improving IR-based Bug Localization with Semantics-Driven Query Reduction

### v1

@@ 4.2 Corpus Indexing
As we aim to localize software bugs using an Information Retrieval (IR)-based solution, it is essential to index all the source documents of a code base (a.k.a., corpus). In our approach, we chose Elasticsearch [42] due to its efficient handling of diverse data types, seamless integration with research tools, and robustness in document indexing. We index all source code (Step 1, Fig. 4) from the 1,578 buggy versions of 42 subject-systems from the Bench4BL dataset.
@@ 4.3 Retrieval of Potentially Buggy Source Documents
We use Elasticsearch to retrieve potentially buggy source documents by executing queries selected from bug reports (Step 2, Fig. 4). Elasticsearch employs its default Standard Analyzer to preprocess queries and retrieve the top-K results (e.g., K = 100) from the corpus. During this retrieval process, we use the default scoring function of Elasticsearch, Okapi BM25 [30], to assess the relevance of the search results. When retrieving suspicious buggy documents, we specify the particular project and version of the bug report where the bug occurred. Unlike other bug localization techniques [19, 18, 11], which consider only the most recent version of a project and may retrieve unrelated documents, our approach ensures that we localize bugs within the correct historical context. This allows us to better mimic real-world scenarios by retrieving documents from the appropriate project and version. At the end of this step, we get a list of potential buggy source documents based on their textual relevance to the bug report.
@@ 4.4 Relevance Estimation using Cross-Encoder
After obtaining results from Elasticsearch, we employ our fine-tuned cross-encoder model to estimate their program semantic relevance to the bug report (Step 3, Fig. 4). Given that the cross-encoder model was trained to differentiate between buggy and bug-free code (Fig. 3.2), its reasoning against the candidate methods should be useful. While source code vocabulary is important, we thus also analyze code-level contexts leveraging the cross-encoder to support bug localization.
We analyze each retrieved source code document by parsing its Abstract Syntax Tree (AST) and extracting its individual methods (Step 1a-1b, 3b). Each method is then paired with the bug report (Step 1c, 3b) and passed into the fine-tuned cross-encoder model (Step 2, 3b), which generates generate a relevance score. This score, ranging from 0 to 1, estimates the likelihood of a method being buggy. This step refines the retrieved documents, enhancing the relevance between the bug report and source code (Step 4, Fig. 4).
@@ 4.5.1 Pre-training of Large Language Models for Software Bugs
We extract salient keywords from a bug report and relevant code segments (Steps 5a, 5b, Fig. 4), using a Transformer-based model. By leveraging the self-attention mechanism [43], Transformer models can focus on the most relevant parts of the text, understand the context, and represent the text comprehensively using high-dimensional vectors. Their representations can be further adapted to different application domains through pre-training using domain specific unstructured data [44].
To help capture salient keywords from bug reports and source code, we pre-trained a transformer model, CodeT5, using bug reports with masking. The process of collecting the pre-training dataset is explained in Section 5.1. We chose the CodeT5 model for embedding generation since it has been trained on both natural language texts and source code, which is ideal for bug reports containing various elements including text, code snippets, and program elements. Masked pre-training [45] involves randomly masking a portion of the input tokens during training, encouraging the model to learn contextual representations by predicting the masked tokens. Bug reports contain valuable information about software issues, including descriptions of bugs, code snippets, and stack traces, etc. By incorporating the masking technique to learn from bug reports, we aimed to enhance CodeT5’s general understanding of the domain-specific language, including its software engineering jargon and bug-related linguistic nuances.
During the pre-training phase, we implemented a masking mechanism where 15% of the tokens in each input sequence (i.e., bug reports) were randomly selected and substituted with a special <extra_id_XX> token. This approach aligns with the original T5 [46] model’s convention, where XX serves as a unique identifier assigned to each masked token. The identifier follows a continuous numbering system, ensuring that each masked token in the input sequence receives a distinct number in sequential order (e.g., <extra_id_0>, <extra_id_1>, <extra_id_2>, and so forth). These input sequences with the masked tokens are then fed into the model to predict the original tokens based on contextual cues. By training the model to reconstruct the original tokens from the masked input, it acquires the ability to capture the general contextual understanding of bugs from bug reports.
@@
Keyword extraction involves selecting a subset of words or phrases that capture the main theme of a text document, which can support various subsequent tasks (e.g., Information Retrieval [47]). For extracting keywords from bug reports, we employ EmbedRank [48] due to its simplicity and compatibility with Transformer models using the KeyBERT [49] library. The Algorithm 1 outlines the keyword extraction process. We begin by pre-processing a bug report using standard text pre-processing techniques, such as removing stop words and punctuation marks. Next, we employ our pre-trained CodeT5 model to embed each token of the bug report. Similarly, we apply the technique to embed the entire bug report. Then we calculate cosine similarity [50] to measure the semantic proximity of each token to the bug report. Based on these scores, we select the top-N most similar keywords (Step 5b, Fig. 4) for the subsequent steps. To maximize diversity in the chosen keywords, we employ the Maximal Marginal Relevance (MMR) [51] algorithm and set the MMR parameter $\lambda=0.5$, following the authors’ recommendation. Since the number of keywords affects retrieval, we also determined the optimal value of N through a controlled experiment, as discussed in $RQ_{2}$.
@@ 5 Experiment
We evaluate our approach using the Bench4BL benchmark dataset and three appropriate performance metrics. We conduct our experiments on a cluster computing system equipped with an NVIDIA GPU with 16 GB of vRAM and compare our technique, IQLoc, against four baseline techniques from the literature. Using our experiments, we attempt to answer four research questions as follows:
@@ 5.1 Dataset Construction
We use Bench4BL [23], a benchmark dataset, for our experiments. Table 2 summarizes the Bench4BL dataset. Since the dataset is relatively old, we refine and extend it with more recent bug reports. Table 3 shows our refined and expanded dataset. In total, we spent over 80 hours refining and expanding the dataset. The following sections detail our refinement process, the collection of new bug reports, and the dataset’s split for training and testing in bug localization.
@@ 5.1.1 Refinement of Bench4BL Dataset:
The Bench4BL [23] dataset contains curated bug reports from various Java-based community projects, including Apache, JBoss, and Spring. It captures 10,017 bug reports from 51 subject systems along with their buggy/fixed versions of code and other relevant metadata. Table 2 provides a summary of the benchmark dataset.
From our initial analysis of the benchmark dataset, we noticed that many projects or bug reports lacked appropriate versioning information, making them unsuitable for our experiment. We thus dropped the bug reports without any version details and ensured that the remaining reports could be traced back to both buggy and bug-free versions of the code. During this process, we encountered several challenges, including cases where either the buggy or fixed version was missing in the code repository. We discarded the bug reports with such discrepancies, ultimately retaining a dataset of 5,753 bug reports from the three large community projects (i.e., Apache, JBoss, and Spring).
@@ 5.1.2 Expanding the Bench4BL Dataset:
To collect bug reports from JIRA-based issue tracking systems, we used the JIRA API [52]. We also filtered issues that were of type ⓐ ‘Bug,’ had a status of ⓑ‘Resolved,’ a resolution of ⓒ ‘Fixed,’ and included both ⓓ a buggy version and ⓔ a fixed version (Fig. 5). Additionally, we considered bug reports containing explicit ⓕ Git pull request links to ensure a proper mapping between the buggy and fixed versions, reducing false positives. Once we collected the bug reports from JIRA, we used the GitHub API to extract the corresponding buggy and fixed files from the linked pull requests. Since a pull request can contain multiple commits, we considered only accepted and verified commits. When multiple commits were associated with a pull request, we included all files across these commits as a part of the fix for that bug report. This approach accounts for bug fixes that evolve over time, often involving multiple developers addressing different aspects of the same bug. Additionally, it captures cases where bug reports were closed and later reopened for further fixes. By including all relevant files, we ensure a comprehensive representation of the bug-fix process. Finally, we ensured that the versions specified in JIRA issues (e.g., semantic version, 4.8.2) matched the version tags in GitHub (e.g., camel-4.8.2) using regex-based validation, as the version strings in JIRA could differ from those in GitHub.
@@
To gain further insight, we classify bug reports in our experimental dataset based on their content, as was done by existing literature [11]:
@@
To classify them, we use regular expressions adapted from the work of Rahman et al. [11]. Fig. 7 shows the classification of our curated dataset.
@@ 5.1.3 Train and Test Set:
For our experiments, we split the dataset into training, validation, and test sets using a 70:10:20 ratio, following the standard machine learning practice for data partitioning [53, 54]. In our dataset splitting, we adopted two strategies, as was done by earlier studies [55, 56]. The first approach utilized random selection with shuffling, where the dataset was randomly divided into training and test sets. To ensure an unbiased random split, we conducted this process five times, generating five randomly shuffled datasets to facilitate separate experiments. This process delivered 5,238 bug reports for training, 1,497 for testing and 748 for validation on random trials in each of the five sets. The second strategy was a time-wise split. This type of splitting divides data into training and test sets based on their chronological order, simulating real-world scenarios where models are trained on past data and evaluated on future instances. For each subject system, we sorted the bug reports by their submission dates, split them individually into train, validation and test sets, and then combined them with similar splits from other systems. This process resulted in 5,236 bug reports for training, 1,501 for testing and 746 bug reports for validation. Table 4 summarizes our curated dataset for the experiments. Also, Fig. 8 shows how classified bug reports (i.e., BRST, BRPE, BRNL) are distributed across our training, validation and test dataset. Note, for random-split, the distribution shows the average of five random-splits.
@@ 5.1.4 Expanding Training Set for Cross-Encoder:
To fine-tune our cross-encoder model, we extract both the buggy and bug-free versions of the code associated with each bug report from the GitHub repository of the respective subject system. Using a Diff tool [57], we compare between buggy and bug-free code and identify the method bodies in the buggy document that were changed to correct the bug. These altered method bodies are selected as positive instances $(q,d^{+})$ for our model training against each bug report, with a label of 1 assigned to them. This process is illustrated in Fig. 3.
@@
To generate negative samples $(q,d^{-})$, we randomly select method bodies associated with different bug reports for a given report at hand. We only consider source code from different subject systems for these negative cases. This process is reiterated four times to yield sufficient training instances for the model, and they are labeled as 0. The choice of four negative instances against each positive instance is based on a study conducted by Huang et al. [58]. They found no notable distinctions across different numbers of negative samples and also endorsed the 4:1 ratio.
@@ 5.1.5 Dataset for Pre-training Model with Bug Reports:
To pre-train an existing baseline language model (e.g., CodeT5) with domain-specific data, we collected thousands of bug reports from GitHub repositories hosting Java-based projects using the GitHub API. Our selection process involved choosing the top 100 repositories based on their star count and thus ensuring that they remained active up to the date of selection. To maintain data integrity, we targeted the issue reports labeled as ’bugs’ and collected reports submitted before April 2024. Additionally, we confirmed that none of these repositories were already included in the Bench4BL dataset to prevent any bias. After collecting the bug reports, we meticulously cleaned them to retain only their textual content using Beautiful Soup [59] and excluded repositories with fewer than five bug reports. Finally, all these steps resulted in a dataset of 70,884 bug reports from 74 repositories. Then, these bug reports were used to pre-train the CodeT5 model.
@@ 5.2 Evaluation Metrics
To evaluate IQLoc in bug localization, we use three widely used metrics- MAP, MRR, and HIT@K. These metrics have been frequently used by the relevant literature on IR-based bug localization [11, 10], and thus are highly relevant to our approach. To perform ablation study involving our cross-encoder model, we used four commonly used metrics – Accuracy, Precision, Recall, and F1 score.
@@
HIT@K or Recall@Top-K [19] refers to the proportion of queries for which a technique returns at least one relevant document among the top $K$ retrieved results. The higher the HIT@K values, the better the performance of a bug localization technique.
@@ 5.3 Evaluting IQLoc
Answering RQ1: Performance of IQLoc
Table 7 presents the performance of IQLoc in terms of Mean Average Precision (MAP), Mean Reciprocal Rank (MRR), and accuracy at different top-K results (HIT@1, HIT@5, HIT@10). For the randomly split datasets, we run the experiments five times, and the table shows the mean performance from five separate runs. On the other hand, we run our experiment once for the time-wise split dataset.
@@
In our approach, we capture the top 100 results retrieved by the Elasticsearch module for subsequent reranking (Step 2, Fig. 4). Previous studies in information retrieval have also used a subset of results to rerank [60, 61, 62]. Our decision was made after carefully analyzing different top-K results. As demonstrated in Table 8, for the random split test set, we observed a HIT@100 of 0.889 and a HIT@200 of 0.929. That is, for $\approx$89% of the bug reports, a relevant result can be found within the top 100 search results, and for $\approx$93% of the reports within the top 200 search results. This is a slight 4.5% increase by considering an additional 100 results. For the time-wise split, this difference is even smaller, only 2.6%. Besides, our reranking step relies on a Transformer-based cross-encoder model, which demands significant computing power. By considering only top 100 results, we thus strike a balance between the relevance of the results and the management of our computational resources.
@@
We also evaluate IQLoc’s performance in localizing different types of bug reports: bug reports containing Stack Trace (ST), Program Element (PE), and Natural Language (NL), as discussed in Section 5.1. In this case, we consider the Elasticsearch as a traditional baseline adapted from Apache Lucene [63]. Table 9 presents the results. In the time-wise split test set (Table 9a), IQLoc outperforms Elasticsearch by 15.77% and 17.15% in MAP and MRR, respectively for bug reports containing stack traces (ST). It also improves such localization by detecting at least one buggy document in the top-10 positions, with gains of 6.41%–20% in HIT@1, HIT@5, and HIT@10. For bug reports with program elements (PE), performance improvements range from 5.6% to 33.46% in the time-wise split, with a 1.2% drop in HIT@10. In contrast, our techniques improves in all metrics for the bug reports containing only natural language (NL), but the gains are smaller. IQLoc achieves only a 4.07% increase in MAP and 4.28% in MRR. HIT@K improvements range from 1.57% to 9.14%, which is lower than the gains for ST and PE. For the random split dataset, we observe a similar trend in localizing bug reports containing stack traces (ST) and program elements (PE). For ST, IQLoc outperforms Elasticsearch by 4.72%–14.02% across all metrics. For PE, the improvement ranges from 1.58% to 12.54%. However, similar to the time-wise split, performance gains for NL bug reports remain minimal, ranging from 0.44% to 2.13% across all metrics except HIT@1.
@@
Answering RQ2: Impact of Query Length and Embedding Models on IQLoc’s Performance

### v2

@@ 4.2 Corpus Indexing
As we aim to localize software bugs using an Information Retrieval (IR)-based solution, it is essential to index all the source documents of a code base (a.k.a., corpus). In our approach, we chose Elasticsearch [elasticsearch] due to its efficient handling of diverse data types, seamless integration with research tools, and robustness in document indexing. We index all source code (Step 1, Fig. 4) from the 1,578 buggy versions of 42 subject-systems from the Bench4BL dataset.
@@ 4.3 Retrieval of Potentially Buggy Source Documents
We use Elasticsearch to retrieve potentially buggy source documents by executing queries selected from bug reports (Step 2, Fig. 4). Elasticsearch employs its default Standard Analyzer to preprocess queries and retrieve the top-K results (e.g., K = 100) from the corpus. During this retrieval process, we use the default scoring function of Elasticsearch, Okapi BM25 [robertson1995okapi_bm25], to assess the relevance of the search results. When retrieving suspicious buggy documents, we specify the particular project and version of the bug report where the bug occurred. Unlike other bug localization techniques [saha_bleuir, ir_localization_lda_buglocator, blizzard], which consider only the most recent version of a project and may retrieve unrelated documents, our approach ensures that we localize bugs within the correct historical context. This allows us to better mimic real-world scenarios by retrieving documents from the appropriate project and version. At the end of this step, we get a list of potential buggy source documents based on their textual relevance to the bug report.
@@ 4.4 Relevance Estimation using Cross-Encoder
After obtaining results from Elasticsearch, we employ our fine-tuned cross-encoder model to estimate their program semantic relevance to the bug report (Step 3, Fig. 4). Given that we trained the cross-encoder model to differentiate between buggy and bug-free code (Fig. 3.2), its reasoning against the candidate methods should be useful. While source code vocabulary is important, our objective is to leverage the cross-encoder’s joint encoding capability (Section 3.2) and capture deep semantic interactions between a bug report and a candidate code through an attention mechanism. Unlike independently learned representations (e.g., embedding-based similarity [yang_word_embedding, polyencoder]), our approach enables us to capture program semantics based on fine-grained, token-level interactions.
We analyze each retrieved source code document by parsing its Abstract Syntax Tree (AST) and extracting its individual methods (Step 1a-1b, 3(b)). Each method is then paired with the bug report (Step 1c, 3(b)) and passed into the fine-tuned cross-encoder model (Step 2, 3(b)), which generates generate a relevance score. This score, ranging from 0 to 1, estimates the likelihood of a method being buggy. This step refines the retrieved documents, enhancing the relevance between the bug report and source code (Step 4, Fig. 4).
@@ 4.5.1 Pre-training of Large Language Models for Software Bugs
We extract salient keywords from a bug report and relevant code segments (Steps 5a, 5b, Fig. 4), using a Transformer-based model. By leveraging the self-attention mechanism [attention_is], Transformer models can focus on the most relevant parts of the text, understand the context, and represent the text comprehensively using high-dimensional vectors. Their representations can be further adapted to different application domains through pre-training using domain specific unstructured data [sun2020sifrank].
To help capture salient keywords from bug reports and source code, we pre-trained a transformer model, CodeT5, using bug reports with masking. The process of collecting the pre-training dataset is explained in Section 5.1. We chose the CodeT5 model for embedding generation since it has been trained on both natural language texts and source code, which is ideal for bug reports containing various elements including text, code snippets, and program elements. Masked pre-training [BERT] involves randomly masking a portion of the input tokens during training, encouraging the model to learn contextual representations by predicting the masked tokens. Bug reports contain valuable information about software issues, including descriptions of bugs, code snippets, and stack traces, etc. By incorporating the masking technique to learn from bug reports, we aimed to enhance CodeT5’s general understanding of the domain-specific language, including its software engineering jargon and bug-related linguistic nuances.
During the pre-training phase, we implemented a masking mechanism where 15% of the tokens in each input sequence (i.e., bug reports) were randomly selected and substituted with a special <extra_id_XX> token. This approach aligns with the original T5 [T5] model’s convention, where XX serves as a unique identifier assigned to each masked token. The identifier follows a continuous numbering system, ensuring that each masked token in the input sequence receives a distinct number in sequential order (e.g., <extra_id_0>, <extra_id_1>, <extra_id_2>, and so forth). These input sequences with the masked tokens are then fed into the model to predict the original tokens based on contextual cues. By training the model to reconstruct the original tokens from the masked input, it acquires the ability to capture the general contextual understanding of bugs from bug reports.
@@
Keyword extraction involves selecting a subset of words or phrases that capture the main theme of a text document, which can support various subsequent tasks (e.g., Information Retrieval [keyword_IR]). For extracting keywords from bug reports, we employ EmbedRank [embedRank] due to its simplicity and compatibility with Transformer models using the KeyBERT [keybert_library] library. The Algorithm 1 outlines the keyword extraction process. We begin by pre-processing a bug report using standard text pre-processing techniques, such as removing stop words and punctuation marks. Next, we employ our pre-trained CodeT5 model to embed each token of the bug report. Similarly, we apply the technique to embed the entire bug report. Then we calculate cosine similarity [cosine_similarity] to measure the semantic proximity of each token to the bug report. Based on these scores, we select the top-N most similar keywords (Step 5b, Fig. 4) for the subsequent steps. To maximize diversity in the chosen keywords, we employ the Maximal Marginal Relevance (MMR) [mmr_algorithm] algorithm and set the MMR parameter $\lambda=0.5$, following the authors’ recommendation. Since the number of keywords affects retrieval, we also determined the optimal value of N through a controlled experiment, as discussed in $RQ_{2}$.
@@ 5 Experiment
We evaluate our approach using the Bench4BL benchmark dataset and three appropriate performance metrics. We conduct our experiments on a cluster computing system equipped with an NVIDIA GPU with 16 GB of vRAM and compare our technique, IQLoc, against eight baseline techniques from the literature. Using our experiments, we attempt to answer four research questions as follows:
@@ 5.1 Dataset Construction
We use Bench4BL [bench4bl], a benchmark dataset, for our experiments. Table 2 summarizes the Bench4BL dataset. Since the dataset is relatively old, we refine and extend it with more recent bug reports. Table 3 shows our refined and expanded dataset. In total, we spent over 80 hours refining and expanding the dataset. The following sections detail our refinement process, the collection of new bug reports, and the dataset’s split for training and testing in bug localization.
@@ 5.1.1 Refinement of Bench4BL Dataset:
The Bench4BL [bench4bl] dataset contains curated bug reports from various Java-based community projects, including Apache, JBoss, and Spring. It captures 10,017 bug reports from 51 subject systems along with their buggy/fixed versions of code and other relevant metadata. Table 2 provides a summary of the benchmark dataset.
From our initial analysis of the benchmark dataset, we noticed that many projects or bug reports lacked appropriate versioning information, making them unsuitable for our experiment. We thus dropped the bug reports without any version details and ensured that the remaining reports could be traced back to both buggy and bug-free versions of the code. During this process, we encountered several challenges, including cases where either the buggy or fixed version was missing in the code repository, as well as instances where the mentioned ground-truth document did not exist. We discarded the bug reports with such discrepancies, ultimately retaining a dataset of 5,753 bug reports from the three large community projects (i.e., Apache, JBoss, and Spring).
@@ 5.1.2 Expanding the Bench4BL Dataset:
To collect bug reports from JIRA-based issue tracking systems, we used the JIRA API [JiraAPI]. We also filtered issues that were of type ⓐ ‘Bug,’ had a status of ⓑ‘Resolved,’ a resolution of ⓒ ‘Fixed,’ and included both ⓓ a buggy version and ⓔ a fixed version (Fig. 5). Additionally, we considered bug reports containing explicit ⓕ Git pull request links to ensure a proper mapping between the buggy and fixed versions, reducing false positives. Once we collected the bug reports from JIRA, we used the GitHub API to extract the corresponding buggy and fixed files from the linked pull requests. Since a pull request can contain multiple commits, we considered only accepted and verified commits. When multiple commits were associated with a pull request, we included all files across these commits as a part of the fix for that bug report. This approach accounts for bug fixes that evolve over time, often involving multiple developers addressing different aspects of the same bug. Additionally, it captures cases where bug reports were closed and later reopened for further fixes. By including all relevant files, we ensure a comprehensive representation of the bug-fix process. Finally, we ensured that the versions specified in JIRA issues (e.g., semantic version, 4.8.2) matched the version tags in GitHub (e.g., camel-4.8.2) using regex-based validation, as the version strings in JIRA could differ from those in GitHub.
@@
To gain further insight, we classify bug reports in our experimental dataset based on their content, as was done by existing literature [blizzard]:
@@
To classify them, we use regular expressions adapted from the work of Rahman et al. [blizzard]. Fig. 7 shows the classification of our curated dataset.
@@ 5.1.3 Train and Test Set:
For our experiments, we split the dataset into training, validation, and test sets using a 70:10:20 ratio, following the standard machine learning practice for data partitioning [dataset_splits, data_splitting]. In our dataset splitting, we adopted two strategies, as was done by earlier studies [time_split, random_split]. The first approach utilized random selection with shuffling, where the dataset was randomly divided into training and test sets. To ensure an unbiased random split, we conducted this process five times, generating five randomly shuffled datasets to facilitate separate experiments. This process delivered 5,238 bug reports for training, 1,497 for testing and 748 for validation on random trials in each of the five sets. The second strategy was a time-wise split. This type of splitting divides data into training and test sets based on their chronological order, simulating real-world scenarios where models are trained on past data and evaluated on future instances. For each subject system, we sorted the bug reports by their submission dates, split them individually into train, validation and test sets, and then combined them with similar splits from other systems. This process resulted in 5,236 bug reports for training, 1,501 for testing and 746 bug reports for validation. Table 4 summarizes our curated dataset for the experiments. Also, Fig. 8 shows how classified bug reports (i.e., BRST, BRPE, BRNL) are distributed across our training, validation and test dataset. Note, for random-split, the distribution shows the average of five random-splits.
@@ 5.1.4 Expanding Training Set for Cross-Encoder:
To fine-tune our cross-encoder model, we extract both the buggy and bug-free versions of the code associated with each bug report from the GitHub repository of the respective subject system. Using a Diff tool [pypiDifftool], we compare between buggy and bug-free code and identify the method bodies in the buggy document that were changed to correct the bug. These altered method bodies are selected as positive instances $(q,d^{+})$ for our model training against each bug report, with a label of 1 assigned to them. This process is illustrated in Fig. 3.
@@
To generate negative samples $(q,d^{-})$, we randomly select method bodies associated with different bug reports for a given report at hand. We only consider source code from different subject systems for these negative cases. This process is reiterated four times to yield sufficient training instances for the model, and they are labeled as 0. The choice of four negative instances against each positive instance is based on a study conducted by Huang et al. [sem_srch_pos_neg_sample]. They found no notable distinctions across different numbers of negative samples and also endorsed the 4:1 ratio.
@@ 5.1.5 Dataset for Pre-training Model with Bug Reports:
To pre-train an existing baseline language model (e.g., CodeT5) with domain-specific data, we collected thousands of bug reports from GitHub repositories hosting Java-based projects using the GitHub API. Our selection process involved choosing the top 100 repositories based on their star count and thus ensuring that they remained active up to the date of selection. To maintain data integrity, we targeted the issue reports labeled as ’bugs’ and collected reports submitted before April 2024. Additionally, we confirmed that none of these repositories were already included in the Bench4BL dataset to prevent any bias. After collecting the bug reports, we meticulously cleaned them to retain only their textual content using Beautiful Soup [BeautifulSoupDocs] and excluded repositories with fewer than five bug reports. Finally, all these steps resulted in a dataset of 70,884 bug reports from 74 repositories. Then, these bug reports were used to pre-train the CodeT5 model.
@@ 5.2 Evaluation Metrics
To evaluate IQLoc in bug localization, we use three widely used metrics- MAP, MRR, and HIT@K. These metrics have been frequently used by the relevant literature on IR-based bug localization [blizzard, rahman2021forgotten], and thus are highly relevant to our approach. To perform ablation study involving our cross-encoder model, we used four commonly used metrics – Accuracy, Precision, Recall, and F1 score.
@@
HIT@K or Recall@Top-K [saha_bleuir] refers to the proportion of queries for which a technique returns at least one relevant document among the top $K$ retrieved results. The higher the HIT@K values, the better the performance of a bug localization technique.
@@ 5.3 Evaluting IQLoc
5.3.1 Answering RQ1: Performance of IQLoc
Table 7 presents the performance of IQLoc in terms of Mean Average Precision (MAP), Mean Reciprocal Rank (MRR), and accuracy at different top-K results (HIT@1, HIT@5, HIT@10). For the randomly split dataset, we repeated our experiments five times with different random seeds, and report the mean performance across these runs. This repeated experiment helps reduce bias from any single random split of the dataset. On the other hand, we run our experiment once for the time-wise split dataset.
@@ 5.3.1 Answering RQ1: Performance of IQLoc
In our approach, we capture the top 100 results retrieved by the Elasticsearch module for subsequent reranking (Step 2, Fig. 4). Previous studies in information retrieval have also used a subset of results to rerank [bugcatcher_multilevel, CE_example_ranking1, CE_example_ranking2]. Our decision was made after carefully analyzing different top-K results. As demonstrated in Table 8, for the random split test set, we observed a HIT@100 of 0.889 and a HIT@200 of 0.929. That is, for $\approx$89% of the bug reports, a relevant result can be found within the top 100 search results, and for $\approx$93% of the reports within the top 200 search results. This is a slight 4.5% increase by considering an additional 100 results. For the time-wise split, this difference is even smaller, only 2.6%. Besides, our reranking step relies on a Transformer-based cross-encoder model, which demands significant computing power. By considering only top 100 results, we thus strike a balance between the relevance of the results and the management of our computational resources.
@@
We also evaluate IQLoc’s performance in localizing different types of bug reports: bug reports containing Stack Trace (ST), Program Element (PE), and Natural Language (NL), as discussed in Section 5.1. In this case, we consider the Elasticsearch as a traditional baseline adapted from Apache Lucene [lucene]. Table 9 presents the results. In the time-wise split test set (Table 9(a)), IQLoc outperforms Elasticsearch by 15.77% and 17.15% in MAP and MRR, respectively for bug reports containing stack traces (ST). It also improves such localization by detecting at least one buggy document in the top-10 positions, with gains of 6.41%–20% in HIT@1, HIT@5, and HIT@10. For bug reports with program elements (PE), performance improvements range from 5.6% to 33.46% in the time-wise split, with a 1.2% drop in HIT@10. In contrast, our techniques improves in all metrics for the bug reports containing only natural language (NL), but the gains are smaller. IQLoc achieves only a 4.07% increase in MAP and 4.28% in MRR. HIT@K improvements range from 1.57% to 9.14%, which is lower than the gains for ST and PE. For the random split dataset, we observe a similar trend in localizing bug reports containing stack traces (ST) and program elements (PE). For ST, IQLoc outperforms Elasticsearch by 4.72%–14.02% across all metrics. For PE, the improvement ranges from 1.58% to 12.54%. However, similar to the time-wise split, performance gains for NL bug reports remain minimal, ranging from 0.44% to 2.13% across all metrics except HIT@1.
@@
5.3.2 Answering RQ2: Impact of Query Length and Embedding Models on IQLoc’s Performance